OpenStudy (anonymous):

The equation below shows the area of a trapezoid, A, with a height of 9 cm, and one base 35 cm. : A = 9 over 2(b + 35) Which of the following formulas correctly solves for the other base, b? b = 2A over 9 + 35 b = 2 multiplied by A over 9 - 35 b = 2 multiplied by A plus 35, all over 9 b = 2 multiplied by A minus 35, all over 9

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

@wio

OpenStudy (zzr0ck3r):

\(A=\dfrac{9}{2(b+35)}\) ?

OpenStudy (anonymous):

A=9/2 (b+35)

OpenStudy (zzr0ck3r):

\[A=\dfrac{9}{2}(b+35)\]?

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

multipply both sides by \(\dfrac{2}{9}\) and what do you have?

OpenStudy (zzr0ck3r):

multiply*

OpenStudy (anonymous):

9/2b+ 315/1

OpenStudy (anonymous):

2/9b

OpenStudy (zzr0ck3r):

\((\dfrac{2}{9})A=(\dfrac{\cancel{2}}{\cancel{9}})\dfrac{\cancel{9}}{\cancel{2}}(b+35)\\\dfrac{2}{9}A=b+35\)

OpenStudy (zzr0ck3r):

you with me?

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

subtract 35 from both sides

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

Hello

OpenStudy (usukidoll):

so leaving off with what zzrocker said we are solving for b and the last step he did was \[(\dfrac{2}{9})A=(\dfrac{\cancel{2}}{\cancel{9}})\dfrac{\cancel{9}}{\cancel{2}}(b+35)\\\dfrac{2}{9}A=b+35 \]

OpenStudy (usukidoll):

there's not that much left to do with this problem... if we want b by itself, what do we have to do?

OpenStudy (anonymous):

-35

OpenStudy (anonymous):

from each side

OpenStudy (usukidoll):

yes

OpenStudy (usukidoll):

\[\frac{2}{9}A - 35 = b\] that's it :)

OpenStudy (usukidoll):

you're looking for this selection right? b = 2 multiplied by A over 9 - 35

OpenStudy (anonymous):

thanks @UsukiDoll

OpenStudy (usukidoll):

can I get a medal? :)

OpenStudy (anonymous):

i just did

OpenStudy (usukidoll):

thank you :)